On March 1, 2017, Bonita Construction Company contracted to construct a factory building for Fabrik Manufacturing Inc. for a total contract price of $8,450,000. The building was completed by October 31, 2019. The annual contract costs incurred, estimated costs to complete the contract, and accumulated billings to Fabrik for 2017, 2018, and 2019 are given below: 2017 2018 2019 Contract costs incurred during the year $2,824,800 $2,086,300 $2,418,900 Estimated costs to complete the contract at 12/31 3,595,200 2,418,900 –0– Billings to Fabrik during the year 3,230,000 3,510,000 1,710,000 (a) Using the percentage-of-completion method, prepare schedules to compute the profit or loss to be recognized as a result of this contract for the years ended December 31, 2017, 2018, and 2019. (Ignore income taxes.) (If answer is 0, please enter 0. Do not leave any fields blank.) (b) Using the completed-contract method, prepare schedules to compute the profit or loss to be recognized as a result of this contract for the years ended December 31, 2017, 2018, and 2019. (Ignore income taxes.) (If answer is 0, please enter 0. Do not leave any fields blank. Enter loss using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)
Answer #1

Part 1

2017

2018

2019

Revenue recognized

3718000

1943500

2788500

Gross profit (loss)

893200

(142800)

369600

2017

2018

2019

Contract price

8,450,000

8,450,000

8,450,000

Actual costs to date

2824800

4911100 (2824800+2086300)

7330000 (4911100+2418900)

Estimated costs to complete

3595200

2418900

0

Total estimated costs

6420000

7330000

7330000

Revenue Recognition:

2017 = 2824800/6420000 =44%*8450000 = 3718000

2018 = 4911100/7330000 = (67%*8450000)-3718000 = 1943500

2019 =8450000-3718000-1943500 =2788500

Gross profit (loss) recognition:

2018 = 3718000-2824800 = 893200

2019=1943500-2086300 = (142800)

2020= 2788500-2418900 = 369600

Part 2

2017

2018

2019

Revenue recognized

0

0

8450000

Gross profit (loss)

0

0

1120000